Triangular Co. will pay a dividend of $3.25 per share, today. The company will continue to pay dividends on a yearly basis, and increase its dividend by 12 percent per year for the next 5 years. After that, the expected dividend growth rate will be negative 3 percent per year forever. If the required return on Triangular stock is 8 percent, and a stock’s share price is equal to the PV of all remaining dividend payments, what will a share of Triangular Co. sell for immediately before the dividend is paid? What will the price of the stock be right after the dividend is paid?
Answer #1

Current Dividend = $ 3.25 = D0, Initial Growth Rate = 12%, Initial Growth Rate Tenure = 5 years

Perpetual Growth Rate = - 3 %

Required Return on Stock = 8%

D1 = 3.25 x 1.12 = $ 3.64

D2 = 3.64 x 1.12 = $ 4.0768

D3 = 4.0768 x 1.12 = $ 4.566

D4 = 4.566 x 1.12 = $ 5.1139

D5 = 5.1139 x 1.12 = $ 5.7276

D6 = 5.7276 x (1-0.03) = $ 5.5556

Horizon Value of Perpetual Dividends = 5.5556 / [0.08 - (-0.03)] = $ 50.507

Present Value of Horizon Value = 50.507 / (1.08)^(5) = $ 34.374

Present Value of Initial Dividends = 3.64 / (1.08) + 4.0768 / (1.08)^(2) + 4.566 / (1.08)^(3) + 5.1139 / (1.08)^(4) + 5.7276 / (1.08)^(5) = $ 18.147

Current Intrinsic Stock Price = PV of Remaining Dividends = 18.147 + 34.374 = $ 52.521

Stock Price Before Dividend Payment = 52.521 + 3.25 = $55.77

Stock Price After Dividend Payment = $ 52.52